if a teacher is 37 years old and one of her students is 11 years old, when will she be twice as old as her student?

Answer by stanbon(75887)

if a teacher is 37 years old and one of her students is 11 years old, when will she be twice as old as her student?
=======
Let the answer be "x"::
------
37 + x = 2(11+x)
---
37 + x = 22 + 2x
x = 15 years
------------------
Cheers,
Stan H.
----------------